OpenStudy (anonymous):

-3√20 mulitipled by √15

OpenStudy (swissgirl):

\( -3 \sqrt(20) *\sqrt(15)\) Ok now the rule goes as follows: \( \sqrt{a}*\sqrt{b}=\sqrt{a*b}\) and \(c\sqrt{a}*\sqrt{b}=c\sqrt{a*b}\)

OpenStudy (swissgirl):

Can you apply these rules to your problem?

OpenStudy (anonymous):

ok ill try and can u tell me if i have found the right answer?

OpenStudy (swissgirl):

Yes definitely :)

OpenStudy (anonymous):

is it -3√300 ?

OpenStudy (swissgirl):

Yessssss :)

OpenStudy (anonymous):

now what am i supposed to do after that to solve?

OpenStudy (anonymous):

find the multiples of 300 ?

OpenStudy (swissgirl):

Okkk now we have to break it down For example: \(\sqrt{28}=\sqrt{7*4}=\sqrt{7*2*2}=2\sqrt{7}\) Yaaaaaaaaaa

OpenStudy (anonymous):

so gcf and then u square u it?

OpenStudy (swissgirl):

Ummm lets first find factors of 300 The way I would begin is 30*10=300 We know that 30=5*6 and 10=5*2 so 300=5*6*5*2

OpenStudy (swissgirl):

Do you follow what I did?

OpenStudy (swissgirl):

hmmmmm 6 can be broken down to 6=3*2

OpenStudy (swissgirl):

Its kind of like we are making a factor tree. Are you familiar with that?

OpenStudy (swissgirl):

So basically 300=3*2*5*5*2

OpenStudy (swissgirl):

hmmmmm